OpenStudy (anonymous):

Solve for x: |x − 2| + 10 = 12 x = 0 and x = 4 x = −4 and x = 0 x = −20 and x = 4 No solution

OpenStudy (lynfran):

|x-2|+10=12 |x-2|=12-10 ....

OpenStudy (campbell_st):

well the 1st step is to subtract 10 from both sides of the equation and you get |x - 2| = 2 since this is an absolute value equation you need to solve x - 2 = 2 and x - 2 = -2 to get the 2 possible solutions